OpenStudy (anonymous):

(-2x+6) ^ 1/5 = (-8+ 10x)^1/5

OpenStudy (anonymous):

\[(-2x+6) ^ {\frac{1}{5}} = (-8+ 10x)^ {\frac{1}{5}} \] Since powers on both sides are eqaul therefore \[(-2x+6) = (-8+ 10x) \rightarrow -2x+6=-8+10x \rightarrow 6+8 =10x+2x\] \[14 =12x \rightarrow x =\frac{14}{12}\] \[x =\frac{7}{6}\]